# Learning functions on symmetric matrices and point clouds via lightweight invariant features

### v2

@@ 3 Invariant functions on symmetric matrices
Theorem 3.1.
There is a closed, $\mathfrak{S}_{n}$-invariant, measure zero set $B\subset\mathcal{S}(n)$ (the “bad set”), such any continuous, $\mathfrak{S}_{n}$-invariant function $\mathcal{S}(n)\setminus B\rightarrow\mathbb{R}$ defined on the complement of $B$ can be uniformly approximated on any compact subset by a multi-layer perceptron that takes the $f$’s as inupt.
@@
The argument follows a standard pattern in machine learning of translating an orbit-separation result (Proposition 3.4 below) into a universal approximation result via the Stone-Weierstrass theorem.
Proof of Theorem 3.1.
@@
where $\mathcal{S}(n)/\mathfrak{S}_{n}$ is the topological quotient of $\mathcal{S}(n)$ by the group action. Since $\mathcal{S}(n)$ is hausdorff and $\mathfrak{S}_{n}$ is compact, the quotient is hausdorff [Bre72, Theorem I.3.1]. Therefore, by the Stone-Weierstrass theorem applied to this quotient, the algebra generated by the functions induced on $\mathcal{S}(n)/\mathfrak{S}_{n}$ by the $f$’s uniformly approximates any continuous function on any compact set on which these functions separate points in the quotient $\mathcal{S}(n)/\mathfrak{S}_{n}$. Pulling back to $\mathcal{S}(n)$, this means that the algebra generated by the $f$’s uniformly approximates any continuous invariant function on any compact set where the $f$’s separate orbits. It is standard [Cyb89, Theorem 1] that a multi-layer perceptron trained on the $f$’s can approximate arbitrary continuous functions in the $f$’s, and in particular, any element of the algebra generated by the $f$’s. Thus, the theorem follows from the following proposition, which will complete the proof. ∎
Proposition 3.4.
@@
The proof uses the main idea from [Thi00, Theorem 11.2]. We relegate it to the Appendix, but discuss the intuition behind the result here (at some length). The $f^{d}$’s and $f^{o}$’s are invariant with respect to an arbitrary permutation of the $X_{ii}$’s and an arbitrary (unrelated) permutation of the $X_{ij}$’s ($i\neq j$); thus, they are invariants of a bigger group action on $\mathcal{S}(n)$ than the action of $\mathfrak{S}_{n}$ described above. Specifically, let $\mathfrak{S}_{n}$ act by arbitrary permutations on the $X_{ii}$’s while fixing the $X_{ij}$’s ($i\neq j$), and let $\mathfrak{S}_{n(n-1)/2}$ act by arbitrary permutations on the $X_{ij}$’s for $i\neq j$ (that respect the relation $X_{ij}=X_{ji}$, i.e., that preserve the symmetry of the matrix $(X_{ij})$) while fixing the $X_{ii}$’s. Since these actions commute with each other (as they are by permutations supported on disjoint sets), they induce an action by the direct product $\Gamma:=\mathfrak{S}_{n}\times\mathfrak{S}_{n(n-1)/2}$ on $\mathcal{S}(n)$. The $f^{d}$’s and $f^{o}$’s are invariant functions for this (bigger) action.
@@
Furthermore, the $f^{d}$’s uniquely characterize the numbers $X_{11},\dots,X_{nn}$ up to order; indeed, one reads off a sorted list of the $X_{ii}$’s by evaluating $f^{d}_{1}(X),\dots,f^{d}_{n}(X)$, by definition of the latter. In a similar way, the $f^{o}$’s uniquely characterize the $\binom{n}{2}$ numbers $X_{ij}$, $1\leq i<j\leq n$ up to order. Thus, taken as a whole, the $f^{d}$’s and $f^{o}$’s yield a complete set of orbit-separating invariants for the action by $\Gamma$ on $\mathcal{S}(n)$ described in the last paragraph.
@@
We would be done (with no need to exclude a “bad set” $B$) if our interest were in this action by $\Gamma=\mathfrak{S}_{n}\times\mathfrak{S}_{n(n-1)/2}$. However, our actual interest is in the action of $\mathfrak{S}_{n}$ given by $\pi X:=P_{\pi}XP_{\pi}^{\top}$ per (1). In the latter action, a permutation $\pi\in\mathfrak{S}_{n}$ acts on both the $X_{ii}$’s and the $X_{ij}$’s at once by permuting the underlying index set over which $i$ and $j$ vary. So what we really want are (complete) invariants for the action of that subgroup $G$ of $\Gamma$ where the entries in a pair $(\sigma,\tau)\in\Gamma=\mathfrak{S}_{n}\times\mathfrak{S}_{n(n-1)/2}$ are induced by the same underlying permutation of the index set. (Note that $G\cong\mathfrak{S}_{n}$, but we introduce the new notation so as to think of it as a subgroup of $\Gamma$.) The $f^{d}$’s and $f^{o}$’s are too invariant: they are invariant under any pair $(\sigma,\tau)$.
This is where
$f^{\star}=\sum_{1\leq i,j\leq n;\;i\neq j}X_{ii}X_{ij}$
@@
enters. For an element $(\sigma,\tau)$ of $\Gamma$ to leave $f^{\star}$ invariant, it must send terms of the form $X_{ii}X_{ij}$ to other such terms. So the pair $(\sigma,\tau)$ of permutations must preserve the “incidence relation” (i.e., the relation of sharing an index) between the $X_{ii}$’s and the $X_{ij}$’s ($i\neq j$). Now $\sigma$ (acting on the $X_{ii}$’s) tells us what the underlying permutation of the indices must be, and the fact that the pair $(\sigma,\tau)$ preserves this incidence relation tells us that $\tau$ must be induced by the same underlying permutation of the indices. (This point is argued carefully in the proof in the Appendix.) Thus the only elements of $\Gamma$ that leave $f^{\star}$ invariant are precisely those belonging to $G$.
General theory then tells us that $f^{\star}$ is able to discriminate between most orbits of $G$ that lie in the same orbit of $\Gamma$. The “bad set” $B$ (where $f^{\star}$ fails to do this) enters because $f^{\star}$’s ability to do this is contingent on the non-vanishing of some polynomials, so $B$ can be taken to be the union of the zero sets of these polynomials.
@@ 4 Invariant functions on point clouds
and similarly for the $\tilde{f}^{d}$’s and $\tilde{f}^{o}$’s (which we refer to collectively as “the $\tilde{f}$’s”). These functions are $\mathrm{O}(d)\times\mathfrak{S}_{n}$-invariant per the discussion above. A priori, because of the inexplicitness of the “bad set” $B$ of the previous section, there is a danger that the set of rank $\leq d$ positive semidefinite matrices lies within it, so we cannot apply Theorem 3.1 or Proposition 3.4 directly in this setting. However, with some additional care, we can prove the following.
Theorem 4.1.
@@
This follows from the following proposition just as Theorem 3.1 followed from Proposition 3.4. ∎
Proposition 4.2.
@@
The proof is in the Appendix. The argument is similar to Proposition 3.4; the main additional ideas are as follows:
@@
  2. 2.
    As with the $f^{d}$’s and $f^{o}$’s of Proposition 3.4, the $\tilde{f}^{d}$’s and $\tilde{f}^{o}$’s identify the Gram matrix of the point cloud up to arbitrary independent permutations of the diagonal and off-diagonal elements. So the argument again comes down to checking that the only ones of these pairs of permutations that fix $\tilde{f}^{*}$ are those that are induced by the same underlying permutation of the original points. But this time, to check this, we need to make sure that all other pairs of permutations not only fail to fix $f^{*}$, but even fail to fix it modulo $I_{d+1}$. So we need to reason a little about the ideal $I_{d+1}$.
@@
5 Reducing the number of features
In this section, we show that for point clouds of $n$ points in dimension $d$ ($d\ll n$), we can reduce the number of invariant features and overall computational complexity from $O(n^{2})$ to $O(n)$, by exploiting that the Gram matrix is low rank. To this end, we make use of two objects: $d$ equivariant identifiers and permutation separators, defined as follows.
@@ 5 Reducing the number of features
Proposition 5.7 below shows that we can take $C_{V}$ to be the matrix with columns the $k$-means centroids of the point cloud $V$ with $k=d$, ordered by the 2-norm. Using $k$-means clustering for quantizing point clouds is a standard practice in data science (e.g.[BCM+20, BGGS21]). In theory, computing the $k$-means centroids is an NP-hard problem [ACKS15], but several randomized algorithms (such as kmeans++ [AV+07]) are very efficient and seem to perform well in practice. However, centroids obtained via randomized algorithms do not exactly fit the theoretical framework we consider here (since we cannot guarantee that the centroids will be a function of the input, let alone $\mathrm{O}(d)$-equivariant or $\mathfrak{S}_{n}$-invariant). There exist deterministic approximation algorithms for the $k$-means centroids that either are or can be modified to be $\mathrm{O}(d)$-equivariant and $\mathfrak{S}_{n}$-invariant, for example the method defined in Theorem 14 of [MVW17]. For simplicity, we prove Proposition 5.7 for the $k$-means optimal centroids (which we may not be able to compute in general), and, also for simplicity, in the numerical experiments in Section 7 we use the standard (randomized) implementation of $k$-means. For the sake of theoretical completeness, we also exhibit an alternative construction for $d$ equivariant identifiers in Proposition 5.8, which are directly computable with $O(n)$ cost (for fixed $d$).
Permutation separators:
@@
A method for constructing permutation separators $D$ as defined above is given in [DG24, Proposition 2.1], with $s=2nd+1$. In practice, we will implement learning models belonging to a function class known as DeepSets [ZKR+17], widely used in machine learning. We describe DeepSets in the next section, and use these models for the numerical experiments in Section 7. Theorems 5 and 6 in [TW24] imply [^3] This implication is not drawn out explicitly in [TW24]. Theorem 5 only states that there exist generically separating DeepSets. However, Theorem 6 cannot hold unless the DeepSets of Theorem 5 are in fact fully separating. that there exist DeepSets with an encoder $\phi$ (see Section 6 for notation) such that the encode-plus-aggregate step $\{x_{i}\}\mapsto\sum_{i}\phi(x_{i})$ serves as a permutation separator with $s=2nd$. This result should be interpreted as saying that DeepSets with $O(n)$ latent dimension are expressive enough to serve the function of the permutation separator. The proof technique requires a full $2nd$ dimensions in the latent space—but as a matter of engineering, for learning a particular target function one may obtain better results using a smaller latent dimension $s$.
@@
It is well-known in invariant theory, in a more general context, that there always exists a separating set of size approximately twice the dimension of the data space [Duf08, Proposition 5.1.1], [Kem24, Theorem 6.1]. However, the proofs involve computing a potentially much larger separating set and projecting it to low dimension, so these theorems do not provide separators computable with $O(n)$ space or time cost. Similarly, [CIMP22, Theorem 12 and Section 3.2.1] provides a methodology for finding twice-dimension-many worst-case separators, but again they are not (at least evidently) computable in $O(n)$ time. The advantage of [TW24, Theorems 5 and 6] for our purposes is that they show that the DeepSets architecture, which is practical for other reasons—including that it allows to uniformly handle point clouds of different sizes—can serve the function of the permutation separators, with not worse than the target computational cost.
Using these ingredients we can prove the following statement.
Proposition 5.4.
@@
We take $\widehat{B}$ to be the union of $B^{\prime}$ and the $\tilde{B}$ of Proposition 4.2. It is clear that $h$ is defined and continuous on the complement of $\widehat{B}$ (and in fact even on the complement of $B^{\prime}$). Similarly, $h$ is $\mathrm{O}(d)\times\mathfrak{S}_{n}$-invariant on $\mathbb{R}^{d\times n}\setminus\widehat{B}$, because given $(U,\pi)\in\mathrm{O}(d)\times\mathfrak{S}_{n}$, we have
@@
Suppose $V_{1},V_{2}\in\mathbb{R}^{d\times n}\setminus\widehat{B}$ lie in different orbits of $\mathrm{O}(d)\times\mathfrak{S}_{n}$. Then they are distinguished by some $\tilde{f}$, by Proposition 4.2 (where the $\tilde{f}$’s are as defined in Section 4). Since the $\tilde{f}$’s can be computed from the Gram matrices $V_{1}^{\top}V_{1}$ and $V_{2}^{\top}V_{2}$, and are invariant with respect to the $\mathfrak{S}_{n}$-conjugation action on these matrices, this implies that $V_{1}^{\top}V_{1}$ and $V_{2}^{\top}V_{2}$ lie in different orbits for the $\mathfrak{S}_{n}$-conjugation action.
@@
Since $D$ is a permutation separator, if $C_{V_{1}}^{\top}V_{1}$ and $C_{V_{2}}^{\top}V_{2}$ differ even up to column permutation (i.e., lie in different orbits for the $\mathfrak{S}_{n}$-action on columns), then this difference is detected by $D$. Thus, either
@@
As in Sections 3 and 4, we immediately get a universal approximation guarantee on the complement of $\widehat{B}$, with the same proof:
Theorem 5.5.
With $h,V$ and $\widehat{B}$ as in Proposition 5.4, any $\mathrm{O}(d)\times\mathfrak{S}_{n}$-invariant function $\mathbb{R}^{d\times n}\setminus\widehat{B}\rightarrow\mathbb{R}$ can be uniformly approximated on any compact subset by a multi-layer perceptron that takes $h(V)$ as input. ∎
@@
In Proposition 5.4, we formed the matrix $C_{V}^{\top}V$, and then processed its columns as a multiset (i.e., in a manner that is invariant to column permutations). This is related to a device used in [HAGD23, Theorem 4], in which $d-1$ points from the original point cloud $V$ are selected arbitrarily, a $d\times d$ matrix $X$ is formed from them in an $\mathrm{O}(d)$-equivariant manner, and the columns of $X^{\top}V^{\prime}$ are then processed as a multiset, where $V^{\prime}$ is $V$ with the selected points omitted. To obtain permutation invariance, the method in [HAGD23] requires to repeat this process over all $\binom{n}{d-1}$ choices of $d-1$ points; the present approach avoids this computational burden by using a single permutation-invariant $C$.
We close the section by verifying the claims made about $d$ equivariant identifiers in Remark 5.1

### v3

@@ 4 Invariant functions on symmetric matrices
Proposition 4.1.
There is a closed, $\mathfrak{S}_{n}$-invariant, measure zero set $B\subset\mathcal{S}(n)$ (the “bad set”), such that any continuous, $\mathfrak{S}_{n}$-invariant function $\mathcal{S}(n)\setminus B\rightarrow\mathbb{R}$ defined on the complement of $B$ can be uniformly approximated on any compact subset by a multi-layer perceptron that takes the $f$’s as input.
@@
The argument follows a standard pattern in machine learning of translating an orbit-separation result (Theorem 4.4 below) into a universal approximation result via the Stone-Weierstrass theorem.
Proof of Proposition 4.1.
@@
where $\mathcal{S}(n)/\mathfrak{S}_{n}$ is the topological quotient of $\mathcal{S}(n)$ by the group action. Since $\mathcal{S}(n)$ is hausdorff and $\mathfrak{S}_{n}$ is compact, the quotient is hausdorff [Bre72, Theorem I.3.1]. Therefore, by the Stone-Weierstrass theorem applied to this quotient, the algebra generated by the functions induced on $\mathcal{S}(n)/\mathfrak{S}_{n}$ by the $f$’s uniformly approximates any continuous function on any compact set on which these functions separate points in the quotient $\mathcal{S}(n)/\mathfrak{S}_{n}$. Pulling back to $\mathcal{S}(n)$, this means that the algebra generated by the $f$’s uniformly approximates any continuous invariant function on any compact set where the $f$’s separate orbits. It is standard [Cyb89, Theorem 1] that a multi-layer perceptron trained on the $f$’s can approximate arbitrary continuous functions in the $f$’s, and in particular, any element of the algebra generated by the $f$’s. Thus, the proposition follows from the following theorem, which will complete the proof. ∎
Theorem 4.4.
@@
We give a proof using Theorem 3.1. A previous version of this paper gave a Galois theory-based proof that is preserved in the Appendix.
Proof.
@@
The proof uses the main idea from [Thi00, Theorem 11.2], and the above Theorem 3.1. The $f^{d}$’s and $f^{o}$’s are invariant with respect to an arbitrary permutation of the $X_{ii}$’s and an arbitrary (unrelated) permutation of the $X_{ij}$’s ($i\neq j$); thus, they are invariants of a bigger group action on $\mathcal{S}(n)$ than the action of $\mathfrak{S}_{n}$ described above. Specifically, let $\mathfrak{S}_{n}$ act by arbitrary permutations on the $X_{ii}$’s while fixing the $X_{ij}$’s ($i\neq j$), and let $\mathfrak{S}_{n(n-1)/2}$ act by arbitrary permutations on the $X_{ij}$’s for $i\neq j$ (that respect the relation $X_{ij}=X_{ji}$, i.e., that preserve the symmetry of the matrix $(X_{ij})$) while fixing the $X_{ii}$’s. Since these actions commute with each other (as they are by permutations supported on disjoint sets), they induce an action by the direct product $\Gamma:=\mathfrak{S}_{n}\times\mathfrak{S}_{n(n-1)/2}$ on $\mathcal{S}(n)$. The $f^{d}$’s and $f^{o}$’s are invariant functions for this (bigger) action.
@@
Furthermore, the $f^{d}$’s uniquely characterize the numbers $X_{11},\dots,X_{nn}$ up to order; similarly, the $f^{o}$’s uniquely characterize the $\binom{n}{2}$ numbers $X_{ij}$, $1\leq i<j\leq n$ up to order. Thus, taken as a whole, the $f^{d}$’s and $f^{o}$’s yield a set of orbit-separating invariants for the action by $\Gamma$ on $\mathcal{S}(n)$ described in the last paragraph.
@@
We would be done (with no need to exclude a “bad set” $B$) if our interest were in this action by $\Gamma=\mathfrak{S}_{n}\times\mathfrak{S}_{n(n-1)/2}$. However, our actual interest is in the action of $\mathfrak{S}_{n}$ given by $\pi X:=P_{\pi}XP_{\pi}^{\top}$ per (1). In the latter action, a permutation $\pi\in\mathfrak{S}_{n}$ acts on both the $X_{ii}$’s and the $X_{ij}$’s at once by permuting the underlying index set over which $i$ and $j$ vary. Let $G$ be the subgroup of $\Gamma$ such that the entries in a pair $(\sigma,\tau)\in\Gamma=\mathfrak{S}_{n}\times\mathfrak{S}_{n(n-1)/2}$ are induced by the same underlying permutation of the index set. (Note that $G\cong\mathfrak{S}_{n}$, but we introduce the new notation so as to think of it as a subgroup of $\Gamma$.)
@@
We now use Theorem 3.1. Let $\mathcal{F}$ be the class of polynomial functions on $\mathcal{S}(n)$. This family forms a vector space that is stable under the action of $\Gamma$, and nonzero polynomials have closed and measure zero vanishing sets. We have $f^{\star}\in\mathcal{F}$; to verify the hypothesis of Theorem 3.1 we need to also check that $\gamma f^{\star}=f^{\star}\Rightarrow\gamma\in G$ for all $\gamma\in\Gamma$. For an element $(\sigma,\tau)$ of $\Gamma$ to leave $f^{\star}$ invariant, it must send terms of the form $X_{ii}X_{ij}$ to other such terms. So the pair $(\sigma,\tau)$ of permutations must preserve the relation of sharing an index between the $X_{ii}$’s and the $X_{ij}$’s ($i\neq j$). Now $\sigma$ (acting on the $X_{ii}$’s) tells us what the underlying permutation of the indices must be, and the fact that the pair $(\sigma,\tau)$ preserves this incidence relation tells us that $\tau$ must be induced by the same underlying permutation of the indices. (This point is argued in more detail in the original proof in the Appendix.) Thus the only elements of $\Gamma$ that leave $f^{\star}$ invariant are precisely those belonging to $G$, and Theorem 3.1 then lets us conclude that the $f^{d}$’s, $f^{o}$’s, and $f^{\star}$ are generically separating for $G$. ∎
@@ 5 Invariant functions on point clouds
and similarly for the $\tilde{f}^{d}$’s and $\tilde{f}^{o}$’s (which we refer to collectively as “the $\tilde{f}$’s”). These functions are $\mathrm{O}(d)\times\mathfrak{S}_{n}$-invariant per the discussion above. A priori, because of the inexplicitness of the “bad set” $B$ of the previous section, there is a danger that the set of rank $\leq d$ positive semidefinite matrices lies within it, so we cannot apply Proposition 4.1 or Theorem 4.4 directly in this setting. However, with some additional care, we can prove the following.
Proposition 5.1.
@@
This follows from the following theorem just as Proposition 4.1 followed from Theorem 4.4. ∎
Theorem 5.2.
@@
The proof is in the Appendix. The argument is similar in spirit and general structure to the proof of Theorem 4.4, but is more involved. The main additional ideas are as follows:
@@
    The action by the bigger group $\Gamma$ of Theorem 4.4 is on $\mathcal{S}(n)$, not on $\mathbb{R}^{d\times n}$. To get around this, we work in an algebraic setting so we can invoke some standard facts about field automorphisms. Thus, in place of Theorem 3.1 we use Galois theory to compute generators for the field of rational invariants of the action of $\mathrm{O}(d)\times\mathfrak{S}_{n}$ on $\mathbb{R}^{d\times n}$ described above—see Corollary A.4—and we use Rosenlicht’s theorem to arrive at generic separation.
  2. 2.
@@
    As with the $f^{d}$’s and $f^{o}$’s of Theorem 4.4, the $\tilde{f}^{d}$’s and $\tilde{f}^{o}$’s identify the Gram matrix of the point cloud up to arbitrary independent permutations of the diagonal and off-diagonal elements. So the argument again comes down to checking that the only ones of these pairs of permutations that fix $\tilde{f}^{*}$ are those that are induced by the same underlying permutation of the original points. But this time, to check this, we need to make sure that all other pairs of permutations not only fail to fix $f^{*}$, but even fail to fix it modulo $I_{d+1}$. So we need to reason a little about the ideal $I_{d+1}$.
  4. 4.
@@
6 Reducing the number of invariants on point clouds
In this section, we show that for point clouds of $n$ points in dimension $d$ ($d\ll n$), we can reduce the number of invariant features and overall computational complexity from $O(n^{2})$ to $O(n)$, by exploiting that the Gram matrix is low rank (see Theorem 6.4). To this end, we make use of two objects: $d$ equivariant identifiers and permutation-orbit separators, defined as follows.
@@ 6 Reducing the number of invariants on point clouds
The $d$ equivariant identifiers is a generic concept that we use as a computational tool in order to reduce the number of invariant features from $O(n^{2})$ to $O(n)$. Standard concepts such as $k$-means centers satisfy this definition even if the data is not clusterable, and can be used as a black-box subroutine. In particular, Proposition 6.7 below shows that we can take $C_{V}$ to be the matrix with columns the $k$-means centroids of the point cloud $V$ with $k=d$, ordered by the 2-norm. Using $k$-means clustering for quantizing point clouds is a standard practice in data science (e.g.[BCM+20, BGGS21]). In theory, computing the $k$-means centroids is an NP-hard problem [ACKS15], but several randomized algorithms (such as kmeans++ [AV+07]) are very efficient and seem to perform well in practice. However, centroids obtained via randomized algorithms do not exactly fit the theoretical framework we consider here (since we cannot guarantee that the centroids will be a function of the input, let alone $\mathrm{O}(d)$-equivariant or $\mathfrak{S}_{n}$-invariant). There exist deterministic approximation algorithms for the $k$-means centroids that either are or can be modified to be $\mathrm{O}(d)$-equivariant and $\mathfrak{S}_{n}$-invariant, for example the method defined in Theorem 14 of [MVW17]. For simplicity, we prove Proposition 6.7 for the $k$-means optimal centroids (which we may not be able to compute in general), and, also for simplicity, in the numerical experiments in Section 8 we use the standard (randomized) implementation of $k$-means. For the sake of theoretical completeness, we also exhibit an alternative construction for $d$ equivariant identifiers in Proposition 6.8, which are directly computable with $O(n)$ cost (for fixed $d$).
Permutation-orbit separators:
@@
A method for constructing permutation-orbit separators $D$ as defined above is given in [DG24, Proposition 2.1], with $s=2nd+1$. In practice, we will implement learning models belonging to a function class known as DeepSets [ZKR+17], widely used in machine learning. We describe DeepSets in the next section, and use these models for the numerical experiments in Section 8. Theorems 5 and 6 in [TW24] imply [^4] This implication is not drawn out explicitly in [TW24]. Theorem 5 only states that there exist generically separating DeepSets. However, Theorem 6 cannot hold unless the DeepSets of Theorem 5 are in fact fully separating. that there exist DeepSets with an encoder $\phi$ (see Section 7 for notation) such that the encode-plus-aggregate step $\{x_{i}\}\mapsto\sum_{i}\phi(x_{i})$ serves as a permutation-orbit separator with $s=2nd$. This result should be interpreted as saying that DeepSets with $O(n)$ latent dimension are expressive enough to serve the function of the permutation-orbit separator. The proof technique requires a full $2nd$ dimensions in the latent space—but as a matter of engineering, for learning a particular target function one may obtain better results using a smaller latent dimension $s$.
@@
It is well-known in invariant theory, in a more general context, that there always exists a separating set of size approximately twice the dimension of the data space [Duf08, Proposition 5.1.1], [Kem24, Theorem 6.1]. However, the proofs involve computing a potentially much larger separating set and projecting it to low dimension, so these theorems do not provide separators computable with $O(n)$ space or time cost. Similarly, [CIMP22, Theorem 12 and Section 3.2.1] provides a methodology for finding twice-dimension-many worst-case separators, but again they are not (at least evidently) computable in $O(n)$ time. The advantage of [TW24, Theorems 5 and 6] for our purposes is that they show that the DeepSets architecture, which is practical for other reasons—including that it allows to uniformly handle point clouds of different sizes—can serve the function of the permutation-orbit separators, with not worse than the target computational cost.
Using these ingredients we provide a construction of $O(n)$ generically separating invariants for $d$-dimensional point clouds with $n$ points. We can prove the following statement.
Theorem 6.4.
@@
We take $\widehat{B}$ to be the union of $B^{\prime}$ and the $\tilde{B}$ of Theorem 5.2. It is clear that $h$ is defined and continuous on the complement of $\widehat{B}$ (and in fact even on the complement of $B^{\prime}$). Similarly, $h$ is $\mathrm{O}(d)\times\mathfrak{S}_{n}$-invariant on $\mathbb{R}^{d\times n}\setminus\widehat{B}$, because given $(U,\pi)\in\mathrm{O}(d)\times\mathfrak{S}_{n}$, we have
@@
Suppose $V_{1},V_{2}\in\mathbb{R}^{d\times n}\setminus\widehat{B}$ lie in different orbits of $\mathrm{O}(d)\times\mathfrak{S}_{n}$. Then they are distinguished by some $\tilde{f}$, by Theorem 5.2 (where the $\tilde{f}$’s are as defined in Section 5). Since the $\tilde{f}$’s can be computed from the Gram matrices $V_{1}^{\top}V_{1}$ and $V_{2}^{\top}V_{2}$, and are invariant with respect to the $\mathfrak{S}_{n}$-conjugation action on these matrices, this implies that $V_{1}^{\top}V_{1}$ and $V_{2}^{\top}V_{2}$ lie in different orbits for the $\mathfrak{S}_{n}$-conjugation action.
@@
Since $D$ is a permutation-orbit separator, if $C_{V_{1}}^{\top}V_{1}$ and $C_{V_{2}}^{\top}V_{2}$ differ even up to column permutation (i.e., lie in different orbits for the $\mathfrak{S}_{n}$-action on columns), then this difference is detected by $D$. Thus, either
@@
As in Sections 4 and 5, we immediately get a universal approximation guarantee on the complement of $\widehat{B}$, with the same proof:
Proposition 6.5.
With $h,V$ and $\widehat{B}$ as in Theorem 6.4, any $\mathrm{O}(d)\times\mathfrak{S}_{n}$-invariant function $\mathbb{R}^{d\times n}\setminus\widehat{B}\rightarrow\mathbb{R}$ can be uniformly approximated on any compact subset by a multi-layer perceptron that takes $h(V)$ as input. ∎
@@
In Theorem 6.4, we formed the matrix $C_{V}^{\top}V$, and then processed its columns as a multiset (i.e., in a manner that is invariant to column permutations). This is related to a device used in [HAGD23, Theorem 4], in which $d-1$ points from the original point cloud $V$ are selected arbitrarily, a $d\times d$ matrix $X$ is formed from them in an $\mathrm{O}(d)$-equivariant manner, and the columns of $X^{\top}V^{\prime}$ are then processed as a multiset, where $V^{\prime}$ is $V$ with the selected points omitted. To obtain permutation invariance, the method in [HAGD23] requires to repeat this process over all $\binom{n}{d-1}$ choices of $d-1$ points; the present approach avoids this computational burden by using a single permutation-invariant $C$.
We close the section with two propositions promised by Remark 6.1. In Proposition 6.7, we verify the claims made in that remark about $k$-means clustering as a choice for the $d$ equivariant identifiers. We note that $k$-means clustering is just a standard method that can be used to produce $d$ equivariant identifiers regardless of whether the data is naturally clusterable. Then, in Proposition 6.8, we provide an alternative implementation of the $d$ equivariant identifiers that does not rely on $k$-means clustering.
@@
Remark 6.9.
Although Theorem 6.4 is formulated in a manner oriented to its purpose in machine learning, we draw out the consequence in the traditional setting of invariant theory, and ask a natural followup question. Because the $C$ of Proposition 6.8 is a polynomial map, and polynomial permutation-orbit separators $D$ of size $O(n)$ do exist as well (for example, the construction in the proof of [TW24, Theorem 5], with appropriate choice of $\ell$), Theorem 6.4 gives a recipe for constructing a family of $O(n)$ polynomials that generically separates the orbits of $\mathrm{O}(d)\times\mathfrak{S}_{n}$ on $\mathbb{R}^{d\times n}$. In the setting of an algebraic group acting on a vector space (or more generally a variety) over an algebraically closed field, generic orbit separation would imply generation of the invariant field, by Rosenlicht’s theorem [Ros56, Theorem 2]. Because $\mathbb{R}$ is not algebraically closed, Theorem 6.4 does not immediately imply a field generation result. But it is natural to ask:
Question 6.10.
For appropriate choice of polynomial permutation-orbit separators $D$, and polynomial $d$ equivariant identifiers $C$, do the entries of $h(V)$ generate the field of rational invariants of $\mathrm{O}(d)\times\mathfrak{S}_{n}$’s action on $\mathbb{R}^{d\times n}$?